OpenStudy (anonymous):

cubed root 200 x^3 y^8

OpenStudy (anonymous):

\[\sqrt[3]{200x ^{3}y ^{8}}\]

OpenStudy (jdoe0001):

 \(\bf \Large {\sqrt[3]{200x^3y^8}\\ \quad \\ \quad \\ \color{green}{200 = 2 \times 2 \times 2\times 5 \times 5\implies 2^3\times 5^2\\ y^8 = y^3\times y^3\times y^2}\\ \quad \\ \sqrt[3]{200x^3y^8}\implies \sqrt[3]{2^3\cdot 5^2x^3y^3y^3y^2}}\)